OpenStudy (anonymous):

If the circumference of a circle is 32 in, find the radius. Find the radius of the circle. Answer 16 in 10.2 in 2.54 in 5.1 in

OpenStudy (anonymous):

what module is this

OpenStudy (anonymous):

multiply 3.14*4

OpenStudy (anonymous):

D

OpenStudy (anonymous):

can i have my medal

OpenStudy (anonymous):

you were wrong

geerky42 (geerky42):

So you can use circumference formula: \(C = 2\pi r\); arrange it a little bit and you have \(\dfrac{C}{2\pi}=r\) plug in value and solve

OpenStudy (anonymous):

I just did that and you are wrong keke

OpenStudy (anonymous):

is it d or not

OpenStudy (anonymous):

Well circumference is 2piR right? So if 2piR = 32 Divide by 2 PiR = 16 And divide both sides by pi...or roughly 3.14 And we get 5.1

OpenStudy (anonymous):

everyone stop typing for a sec

OpenStudy (anonymous):

wolf what module of flvs is this in

OpenStudy (anonymous):

WOLF what module and lesson is this in

OpenStudy (cp9454):

D

OpenStudy (anonymous):

Shut up

OpenStudy (anonymous):

I am asking wolf a question I did this question before but I forgot the answer

OpenStudy (anonymous):

D

OpenStudy (anonymous):

xD